focus (2,5) and directrix y=3
write standard form of the equation of the parabola with the given criteria.

Answer by lwsshak3(11628)

focus (2,5) and directrix y=3
write standard form of the equation of the parabola with the given criteria.
***
Given parabola opens upward. (directrix below focus)
Its basic form of equation: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex.
y-coordinate of vertex=midpoint between focus and directrix on the axis of symmetry=(5+3)/2=4
x-coordinate of vertex=2
vertex: (2,4)
axis of symmetry: x=2
p=1 (distance from vertex to focus or directrix on the axis of symmetry)
4p=4
Equation of given parabola: (x-2)^2=4(y-4)